write an equation in standard form of the hyperpola with foci at (+3, 0) (-3,0) if the difference in the distances from a poin (x,y) on the hyperbola to the foci is 4
Answer by lwsshak3(11628)

write an equation in standard form of the hyperpola with foci at (+3, 0) (-3,0) if the difference in the distances from a poin (x,y) on the hyperbola to the foci is 4.
**
Standard form of equation for a hyperbola with horizontal transverse axis:
(x-h)^2/a^2-(y-k)^2/b^2=1, (h,k)=(x,y) coordinates of center
For given hyperbola:
center:(0,0)
Difference in distances from a point (x,y) on the hyperbola to each of the foci is 4=2a
a=2
a^2=4
c=3
c^2=9
c^2=a^2+b^2
b^2=c^2-a^2=9-4=5
Equation of given hyperbola:
x^2/4-y^2/5=1